# Neutrino mixing parameters and masses from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach

### v1

@@ 4 Typical interesting breaking patterns
By performing a $\chi^{2}$ analysis on the formulated lepton flavor symmetry breaking patterns for the representation assignments $L\sim\bm{3_{0}}$, $\phi_{\text{atm}}\sim\bm{\bar{3}_{0}}$ and $\phi_{\text{sol}}\sim\bm{\bar{3}_{1}}$ within the flavor symmetry $\Delta(96)\rtimes H_{CP}$ under the tri-direct CP approach, we derived that ̵42 symmetry breaking patterns yield predictions agree with the experimental data very well. It is beyond the scope of our study to explore all the viable cases in detail and to present a complete graphical treatment of the predictions for each breaking pattern. In what follows we consider five representative cases of them and present their predictions for the lepton mixing angles, CP violation phases and neutrino masses, in which the quality of the results can be thoroughly appreciated.
@@ 4.1 The breaking pattern $\mathcal{F}_{6}$
For this symmetry breaking pattern, the corresponding residual flavor symmetry combination is given by $(G_{l},G_{atm},G_{sol})=(Z_{4}^{c},Z_{2}^{a^{2}bd^{2}},Z_{2}^{abc})$, with the residual CP transformations of the atmospheric and solar neutrino sectors $(X^{\text{atm}}_{\bm{\bar{3}_{0}}},X^{\text{sol}}_{\bm{\bar{3}_{1}}})=(\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(d^{2}))$. In this case, the diagonalization matrix $U_{l}$ of the charged lepton mass matrix is identity matrix and the VEV alignment of flavor fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ can be obtained from table LABEL:tab:invariant_VEVs, i.e.
@@
where the VEV magnitudes $v_{\phi_{a}}$ and $v_{\phi_{s}}$ are real. Then the low-energy effective light neutrino mass matrix read as
@@
where the parameters $m_{a}$, $r$ and $\eta$ are defined in Eq. (2.10). Note an overall unphysical phase (the phase of $m_{a}$) of $m_{\nu}$ can be omitted and it will be neglected for the other cases in the following. The light neutrino mass matrix $m_{\nu}$ in Eq. (4.2) can be simplified into a block diagonal form $m^{\prime}_{\nu}$ by performing a unitary transformation $U_{\nu 1}$ with $m^{\prime}_{\nu}=U_{\nu 1}^{T}m_{\nu}U_{\nu 1}$, where the unitary matrix $U_{\nu 1}$ and the mass matrix $m^{\prime}_{\nu}$ are
@@
$\sin^{2}\theta_{13}$ $=$ $\frac{2\sin^{2}\theta}{x^{2}-2x+3},\qquad\sin^{2}\theta_{12}=\frac{2\cos^{2}\theta}{x^{2}-2x+2+\cos 2\theta},$
@@
All mixing parameters and neutrino masses are functions of $m_{a}$, $x$, $r$ and $\eta$. Experimental measurements can be matched by choosing suitable values for these parameters. Our numerical analysis of the breaking pattern $\mathcal{F}_{6}$ provides best fit values, summarized in table 3. After scanning the four free parameters of the breaking pattern and enforcing the experimental $3\sigma$ ranges for the mixing angles and Dirac CP phase [50], we present the predicted ranges for neutrino mixing parameters and the absolute neutrino masses in table LABEL:tab:viable_regions.
@@
Figure 4: Contour plots of $\sin^{2}\theta_{23}$, $\sin^{2}\theta_{13}$, $\delta_{CP}$ and $m_{2}^{2}/m_{3}^{2}$ in the $\eta$/$\pi-r$ plane for breaking pattern $\mathcal{F}_{6}$ with $x=3$. On the right panel, we show the predictions for mixing parameters and mass ratio as functions of $\sin\theta_{13}$ for $\mathcal{F}_{6}$ with $x=3$ and $\eta=5\pi/3$.
@@
$\mathcal{F}_{6}$ $\left(\sqrt{\frac{2}{3}},\frac{1}{\sqrt{6}},\frac{1}{\sqrt{6}}\right)^{T}$ $3$ $\frac{5}{3}$ 26.772 0.100 6.421 0.02244 0.318 0.487 1.483 1.599 8.609 50.158 2.688
@@
$\mathcal{F}_{7}$ $\left(\sqrt{\frac{2}{3}},\frac{1}{\sqrt{6}},\frac{1}{\sqrt{6}}\right)^{T}$ $3$ $\frac{4}{3}$ 26.786 0.100 11.598 0.02230 0.318 0.512 1.517 0.402 8.590 50.170 2.681
@@
$\mathcal{F}_{4}$ $\frac{1}{\sqrt{38}}\left(5,2,3\right)^{T}$ $5$ 00 $2.907$ $0.338$ $4.003$ $0.02239$ $0.327$ $0.486$ $1$ 00 $8.653$ $50.131$ $3.889$
@@
$\mathcal{T}_{10}$ $\left(\sqrt{\frac{37}{57}},\sqrt{\frac{10}{57}},\sqrt{\frac{10}{57}}\right)^{T}$ $4$ $1$ $19.289$ $0.0602$ $14.058$ $0.02270$ $0.336$ $0.482$ $1.474$ $1.165$ $8.465$ $50.244$ $3.868$
@@
Table 5: The best fit values of the five interesting breaking patterns $\mathcal{F}_{6}$, $\mathcal{F}_{7}$, $\mathcal{F}_{4}$, $\mathcal{T}_{10}$ and $\mathcal{S}_{2}$ for some benchmark values of parameters $x$ and $\eta$, where $\bm{\hat{v}}_{\text{fix}}$ represents the absolute values of the first of PMNS matrix.
@@ 4.2 The breaking pattern $\mathcal{F}_{7}$
The neutrino mass matrix $m^{\prime}_{\nu}$ can be diagonalized by performing the unitary transformation $U_{\nu 2}$. As a consequence, the lepton mixing matrix is take the following form
@@
Then the analytical expressions for the three lepton mixing angles and CP invariants can be readily derived. When treating $m_{a}$, $x$, $r$ and $\eta$ as free parameters, we present the best fit values of the neutrino masses and mixing parameters in table 3. Following a thorough exploration of the parameter space, we determine the predicted ranges for input parameters and observables, as systematically compiled in table LABEL:tab:viable_regions.
@@
Figure 5: Contour plot of $\delta_{CP}/\pi$ within $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane, and the dimensionless observable quantities $\sin^{2}\theta_{23}$, $\sin^{2}\theta_{13}$, $\delta_{CP}$ and $m_{2}^{2}/m_{3}^{2}$ in the $\eta$/$\pi-r$ plane for the breaking pattern $\mathcal{F}_{7}$ with $x=(2+\sqrt{17})/2\approx 3$.
@@
The left panel of figure 5 illustrates the contour map of $\delta_{CP}/\pi$ in the $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane. With the parameter $x$ held constant, all mixing parameters and the mass ratio $m_{2}/m_{3}$ become fully determined by the two input parameters $r$ and $\eta$. The red shaded region in the left panel represents the viable parameter space, where $r$ and $\eta$ are numerically scanned across their experimentally allowed ranges while satisfying $3\sigma$ constraints on the five dimensionless observable quantities in table 1. This constrained parameter space reveals that both the atmospheric mixing angle $\theta_{23}$ and the Dirac CP violation phase $\delta_{CP}$ are predicted to lie in narrow intervals near their theoretically maximal values. The right panel of figure 5 displays contour bands corresponding to those dimensionless observable quantities up to $\theta_{12}$ within the $\eta/\pi-r$ parameter plane. It show that the input parameters $r$ and $\eta$ are restricted to a narrow region of the parameter space. The solar mixing angle $\theta_{12}$ is governed by the reactor angle $\theta_{13}$ through the sum rule derived in Eq. (4.18), establishing a predictive relationship between these two oscillation parameters.
@@
In the present analysis, we direct our attention to parametrization $x=(2+\sqrt{17})/2$. This specific choice generates a interesting mixing matrix structure characterized by its first column component explicitly given in Eq. (4.16). The phase parameter $\eta$ is constrained to $\eta=4\pi/3$, a value that proves theoretically accessible for implementation within concrete model frameworks. This optimized parameter selection yields a phenomenologically successful and predictive description of lepton mixing parameters and neutrino masses. All lepton mixing parameters and the neutrino mass ratio $m_{2}/m_{3}$ are now determined by a single dimensionless parameter $r$. Upon elimination of the free parameter $r$, sum rules emerge among these quantities. While their explicit analytical forms are omitted here, we present the numerical correlations graphically. Figure 6 illustrates the functional dependence of $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$, the Dirac CP phase $\delta_{CP}$, the Majorana CP phase $\beta$ and mass ratio $m_{2}/m_{3}$ on $\sin\theta_{13}$. Furthermore, we numerically scan over the parameter space of $m_{a}$ and $r$. We find that the input parameters, the neutrino masses and mixing parameters are predicted to lie in the following rather narrow regions
@@
These predictions could be tested by JUNO experiment [91] in combination with DUNE [92] or T2HK [93].
@@ 4.3 The breaking pattern $\mathcal{F}_{4}$
As demonstrated in Eq. (2.19), the nonzero enters $y$, $z$ and $w$ of $m^{\prime}_{\nu}$ take the following form
@@
The lepton mixing angles and CP invariants $J_{CP}$ and $I_{1}$ can be directly extracted through this mixing matrix, though we refrain from providing their explicit mathematical derivations here. By treating $m_{a}$, $x$, $r$ and $\eta$ as free parameters, we determine the best fit neutrino masses and mixing angles, please see table 3. A full parameter scan yields allowed ranges for input parameters, mixing angles, CP phases, and neutrino masses, and we summarizes them in table LABEL:tab:viable_regions.
@@
We now focus on the simple values $x=5$ and $\eta=5\pi/3$, which successfully predicts lepton mixing and neutrino masses. Moreover, the corresponding vacuum alignments of the flavon fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$, and the phase $\eta$ should be easily achievable in a realistic model. Now, all lepton mixing parameters and $m_{2}/m_{3}$ depend solely on the dimensionless parameter $r$, which can be determined by $\theta_{13}$ measurements. In figure 7, we shows correlations between $\theta_{13}$ and other observables. If $m_{2}/m_{3}$ is constrained to its $3\sigma$ range, the input parameters and all mixing parameters fall within narrow predicted regions.
@@
Note that the Dirac CP violation phase $\delta_{CP}$ is predicted to lie in narrow intervals near $\pi$ for the breaking pattern $\mathcal{F}_{4}$ with $x=5$ and $\eta=5\pi/3$. It yield distinct predictions for $\delta_{CP}$ compared to the original Littlest Seesaw model [60] and the breaking patterns $\mathcal{F}_{6}$ and $\mathcal{F}_{7}$.

### v2

@@ 4 Typical interesting breaking patterns
By performing a $\chi^{2}$ analysis on the formulated lepton flavor symmetry breaking patterns for the representation assignments $L\sim\bm{3_{0}}$, $\phi_{\text{atm}}\sim\bm{\bar{3}_{0}}$ and $\phi_{\text{sol}}\sim\bm{\bar{3}_{1}}$ within the flavor symmetry $\Delta(96)\rtimes H_{CP}$ under the tri-direct CP approach, we have derived that 54 symmetry breaking patterns yield predictions that agree with the experimental data very well. It is beyond the scope of our study to explore all the viable cases in detail and to present a complete graphical treatment of the predictions for each breaking pattern. In what follows we consider six (five for NO and one for IO) representative cases of them and present their predictions for the lepton mixing angles, CP violation phases and neutrino masses, in which the quality of the results can be thoroughly appreciated.
@@ 4.1 The breaking pattern $\mathcal{F}_{6}$
For this symmetry breaking pattern, the corresponding residual flavor symmetry combination is given by $(G_{l},G_{atm},G_{sol})=(Z_{4}^{c},Z_{2}^{a^{2}bd^{2}},Z_{2}^{abc})$, with the residual CP transformations of the atmospheric and solar neutrino sectors $(X^{\text{atm}}_{\bm{\bar{3}_{0}}},X^{\text{sol}}_{\bm{\bar{3}_{1}}})=(\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(d^{2}))$. In this case, the diagonalization matrix $U_{l}$ of the charged lepton mass matrix is the identity matrix and the VEV alignment of flavor fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ can be obtained from table LABEL:tab:invariant_VEVs, i.e.
@@
where the VEV magnitudes $v_{\phi_{a}}$ and $v_{\phi_{s}}$ are real. Then the low-energy effective light neutrino mass matrix reads as
@@
where the parameters $m_{a}$, $r$ and $\eta$ are defined in Eq. (2.10). Note that an overall unphysical phase (the phase of $m_{a}$) of $m_{\nu}$ can be omitted and it will be neglected for the other cases in the following. The light neutrino mass matrix $m_{\nu}$ in Eq. (4.2) can be simplified into a block diagonal form $m^{\prime}_{\nu}$ by performing a unitary transformation $U_{\nu 1}$ with $m^{\prime}_{\nu}=U_{\nu 1}^{T}m_{\nu}U_{\nu 1}$, where the unitary matrix $U_{\nu 1}$ and the mass matrix $m^{\prime}_{\nu}$ are
@@
$\sin^{2}\theta_{13}$ $=$ $\frac{2\sin^{2}\theta}{x^{2}-2x+3},\qquad\sin^{2}\theta_{12}=\frac{2\cos^{2}\theta}{x^{2}-2x+2+\cos 2\theta},\$
@@
All mixing parameters and neutrino masses are functions of $m_{a}$, $x$, $r$ and $\eta$. Experimental measurements can be matched by choosing suitable values for these parameters. Our numerical analysis of the breaking pattern $\mathcal{F}_{6}$ provides best fit values, summarized in table LABEL:tab:viable_best_fit. After scanning the four free parameters of the breaking pattern and enforcing the experimental $3\sigma$ ranges for the mixing angles and Dirac CP phase [50], we present the predicted ranges for neutrino mixing parameters and the absolute neutrino masses in table LABEL:tab:viable_regions.
@@
Figure 5: Contour plots of $\sin^{2}\theta_{23}$, $\sin^{2}\theta_{13}$, $\delta_{CP}$ and $m_{2}^{2}/m_{3}^{2}$ in the $\eta$$\pi-r$ plane for breaking pattern $\mathcal{F}_{6}$ with $x=3$. On the right panel, we show the predictions for mixing parameters and mass ratio as functions of $\sin\theta_{13}$ for $\mathcal{F}_{6}$ with $x=3$ and $\eta=5\pi/3$.
@@
$\mathcal{F}_{6}$ $\left(\sqrt{\frac{2}{3}},\frac{1}{\sqrt{6}},\frac{1}{\sqrt{6}}\right)^{T}$ 33 $\frac{5}{3}$ 26.772 0.100 6.421 0.02244 0.318 0.487 1.483 1.599 8.609 50.158 2.688
@@
$\mathcal{F}_{7}$ $\left(\sqrt{\frac{2}{3}},\frac{1}{\sqrt{6}},\frac{1}{\sqrt{6}}\right)^{T}$ 33 $\frac{4}{3}$ 26.786 0.100 11.598 0.02230 0.318 0.512 1.517 0.402 8.590 50.170 2.681
@@
$\mathcal{F}_{4}$ $\frac{1}{\sqrt{38}}\left(5,2,3\right)^{T}$ 55 0 $2.907$ $0.338$ $4.003$ $0.02239$ $0.327$ $0.486$ 11 0 $8.653$ $50.131$ $3.889$
@@
$\mathcal{T}_{10}$ $\left(\sqrt{\frac{37}{57}},\sqrt{\frac{10}{57}},\sqrt{\frac{10}{57}}\right)^{T}$ 44 11 $19.289$ $0.0602$ $14.058$ $0.02270$ $0.336$ $0.482$ $1.474$ $1.165$ $8.465$ $50.244$ $3.868$
@@
For the IO case
$\bm{\hat{v}}_{\text{fix}}$ $x$ $\eta/\pi$ $\frac{m_{a}}{\text{meV}}$ $r$ $\chi^{2}_{\text{min}}$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{12}$ $\sin^{2}\theta_{23}$ $\delta_{CP}/\pi$ $\beta/\pi$ $\frac{m_{1}}{\text{meV}}$ $\frac{m_{2}}{\text{meV}}$ $\frac{m_{\text{ee}}}{\text{meV}}$
$\mathcal{F}_{2}^{\prime}$ $\frac{1}{2\sqrt{11}}\left(1,5,3\sqrt{2}\right)^{T}$ 77 $\frac{52}{27}$ $21.847$ $0.0587$ $14.672$ 0.02273 $0.300$ $0.581$ $1.700$ $1.632$ $49.284$ $49.991$ $41.839$
$\frac{79}{41}$ $21.670$ $0.0587$ 17.345 0.02273 $0.318$ $0.581$ $1.708$ $1.639$ $48.829$ $49.646$ $41.484$
Table 5: The best fit values of the five interesting breaking patterns $\mathcal{F}_{6}$, $\mathcal{F}_{7}$, $\mathcal{F}_{4}$, $\mathcal{T}_{10}$, $\mathcal{S}_{2}$ and $\mathcal{F}_{2}^{\prime}$ for some benchmark values of parameters $x$ and $\eta$, where $\bm{\hat{v}}_{\text{fix}}$ represents the absolute values of the first (third) of PMNS matrix for the NO (IO) case.
@@ 4.2 The breaking pattern $\mathcal{F}_{7}$
The neutrino mass matrix $m^{\prime}_{\nu}$ can be diagonalized by performing the unitary transformation $U_{\nu 2}$. As a consequence, the lepton mixing matrix takes the following form
@@
Then the analytical expressions for the three lepton mixing angles and CP invariants can be readily derived. When treating $m_{a}$, $x$, $r$ and $\eta$ as free parameters, we present the best fit values of the neutrino masses and mixing parameters in table LABEL:tab:viable_best_fit. Following a thorough exploration of the parameter space, we determine the predicted ranges for input parameters and observables, as systematically compiled in table LABEL:tab:viable_regions.
@@
Figure 6: Contour plot of $\delta_{CP}/\pi$ within $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane, and the dimensionless observable quantities $\sin^{2}\theta_{23}$, $\sin^{2}\theta_{13}$, $\delta_{CP}$ and $m_{2}^{2}/m_{3}^{2}$ in the $\eta$$\pi-r$ plane for the breaking pattern $\mathcal{F}_{7}$ with $x=(2+\sqrt{17})/2\approx 3$.
@@
The left panel of figure 6 illustrates the contour map of $\delta_{CP}/\pi$ in the $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane. With the parameter $x$ held constant, all mixing parameters and the mass ratio $m_{2}/m_{3}$ become fully determined by the two input parameters $r$ and $\eta$. The red shaded region in the left panel represents the viable parameter space, where $r$ and $\eta$ are numerically scanned across their experimentally allowed ranges while satisfying $3\sigma$ constraints on the five dimensionless observable quantities in table 1. This constrained parameter space reveals that both the atmospheric mixing angle $\theta_{23}$ and the Dirac CP violation phase $\delta_{CP}$ are predicted to lie in narrow intervals near their theoretically maximal values. The right panel of figure 6 displays contour bands corresponding to those dimensionless observable quantities except $\theta_{12}$ within the $\eta/\pi-r$ parameter plane. It shows that the input parameters $r$ and $\eta$ are restricted to a narrow region of the parameter space. The solar mixing angle $\theta_{12}$ is governed by the reactor angle $\theta_{13}$ through the sum rule derived in Eq. (4.18), establishing a predictive relationship between these two oscillation parameters.
@@
In the present analysis, we direct our attention to the parametrization $x=(2+\sqrt{17})/2$. This specific choice generates an interesting mixing matrix structure characterized by its first column component explicitly given in Eq. (4.16). The phase parameter $\eta$ is constrained to $\eta=4\pi/3$, a value that proves theoretically accessible for implementation within concrete model frameworks. This optimized parameter selection yields a phenomenologically successful and predictive description of lepton mixing parameters and neutrino masses. All lepton mixing parameters and the neutrino mass ratio $m_{2}/m_{3}$ are now determined by a single dimensionless parameter $r$. Upon elimination of the free parameter $r$, sum rules emerge among these quantities. While their explicit analytical forms are omitted here, we present the numerical correlations graphically. Figure 7 illustrates the functional dependence of $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$, the Dirac CP phase $\delta_{CP}$, the Majorana CP phase $\beta$ and mass ratio $m_{2}/m_{3}$ on $\sin\theta_{13}$. Furthermore, we numerically scan over the parameter space of $m_{a}$ and $r$. We find that the input parameters, the neutrino masses and mixing parameters are predicted to lie in the following rather narrow regions
@@
These predictions could be tested by the JUNO experiment [95] in combination with DUNE [96] or T2HK [97].
@@ 4.3 The breaking pattern $\mathcal{F}_{4}$
As demonstrated in Eq. (2.19), the nonzero entries $y$, $z$ and $w$ of $m^{\prime}_{\nu}$ take the following form
@@
The lepton mixing angles and CP invariants $J_{CP}$ and $I_{1}$ can be directly extracted through this mixing matrix, though we refrain from providing their explicit mathematical expressions here. By treating $m_{a}$, $x$, $r$ and $\eta$ as free parameters, we determine the best fit neutrino masses and mixing angles. Please see table LABEL:tab:viable_best_fit. A full parameter scan yields allowed ranges for input parameters, mixing angles, CP phases, and neutrino masses, and we summarize them in table LABEL:tab:viable_regions.
@@
We now focus on the simple values $x=5$ and $\eta=5\pi/3$, which successfully predict lepton mixing and neutrino masses. Moreover, the corresponding vacuum alignments of the flavon fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$, and the phase $\eta$ should be easily achievable in a realistic model. Now, all lepton mixing parameters and $m_{2}/m_{3}$ depend solely on the dimensionless parameter $r$, which can be determined by $\theta_{13}$ measurements. In figure 8, we show correlations between $\theta_{13}$ and other observables. If $m_{2}/m_{3}$ is constrained to its $3\sigma$ range, the input parameters and all mixing parameters fall within narrow predicted regions.
@@
Note that the Dirac CP violation phase $\delta_{CP}$ is predicted to lie in narrow intervals near $\pi$ for the breaking pattern $\mathcal{F}_{4}$ with $x=5$ and $\eta=5\pi/3$. It yields distinct predictions for $\delta_{CP}$ compared to the original Littlest Seesaw model [60] and the breaking patterns $\mathcal{F}_{6}$ and $\mathcal{F}_{7}$.